solve the differential equation dy/dx + y cos x = y^3sin2x


1
Expert's answer
2021-02-24T06:50:51-0500

"\frac{dy}{dx}+y\cos{x}=y^3\sin{2x}"

Solution:

Divide the left and right sides of the equation by "y^3" :

"\frac{1}{y^3}\frac{dy}{dx}+\frac{1}{y^2}\cos{x}=\sin{2x}"

"\frac{1}{y^2}=z" , "z'=-2\frac{y'}{y^3}" , "\frac{y'}{y^3}=-\frac{z'}{2}" .

"-\frac{z'}{2}+z\cos{x}=\sin{2x}"

"z'-2z\cos{x}=-2\sin{2x}"

"z=uv" , "z'=u'v+v'u"

"u'v+v'u-2uv\cos{x}=-2\sin{2x}"

"u'v+u(v'-2v\cos{x})=-2\sin{2x}"

Let's compose and solve the system:

"\begin{cases}
 v'-2v\cos{x}=0 \\
 u'v=-2\sin{2x}
\end{cases}"

From the first equation:

"\frac{dv}{v}=2\cos{x}dx"

"\ln|v|=2\sin{x}"

"v=e^{2\sin{x}}"

Substitute "v" into the second equation:

"u'e^{2\sin{x}}=-2\sin{2x}"

"du=-\frac{2\sin{2x}dx}{e^{2\sin{x}}}"

"\displaystyle u=-\int\frac{4\sin{x}\cos{x}dx}{e^{2\sin{x}}}=-\int\frac{4\sin{x}d\sin{x}}{e^{2\sin{x}}}="

"\displaystyle\frac{2\sin{x}+1}{e^{2\sin{x}}}+C"

"z=uv=e^{2\sin{x}}(\displaystyle\frac{2\sin{x}+1}{e^{2\sin{x}}}+C)="

"\displaystyle2\sin{x}+1+Ce^{2\sin{x}}"

"\frac{1}{y^2}=\displaystyle2\sin{x}+1+Ce^{2\sin{x}}"

Answer: "\frac{1}{y^2}=\displaystyle2\sin{x}+1+Ce^{2\sin{x}}" .